Determine the empirical and molecular formula given the percent compositions or masses.




26% carbon, 3.3% Hydrogen, 70.3%Oxygen Molar mass=91.0g/mol

1
Expert's answer
2021-11-18T12:46:02-0500

M(C) = 12.01 amu

M(H) = 1.0078 amu

M(O) = 15.99 amu

Ratio(C) "= \frac{\%}{M} = \frac{26}{12.01} = 2.16"

Ratio (H) "= \frac{3.3}{1.0078} = 3.27"

Ratio(O) "= \frac{70.3}{15.99} = 4.39"

Simple ratio (C) "= \frac{2.16}{2.16} = 1"

SR(H) "= \frac{3.27}{2.16} = 1.5"

SR(O) "= \frac{4.39}{2.16} = 2.0"

Proportion: 1: 1.5 : 2.0 = 2 : 3 : 4

Empirical formula: C2H3O4

M(C2H3O4) =

"= 2 \times 12.01 + 3 \times 1.0078 + 4 \times 15.99 \\
= 24.02 + 3.0234 + 63.96 \\
= 91.00 \;g/mol"

Molecular formula: C2H3O4